Factor completely
9x^2 + 24xy - 9y^2
I know the GCF is 3, but I don't know what to do with the two different variables. Thank you

Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

9X^2+24XY-9Y^2
(9X-3)(X+3)
3(3X-1)(X+3)

Answer by stanbon(75887)

Factor completely
9x^2 + 24xy - 9y^2
=3(3x^2+8xy-3y^2)
Think of two numbers whose product is 3*-3=-9 and whose sum is 8
The numbers are +9 and -1
So rewrite the problem putting 9xy-1xy in place of 8xy
=3(3x^2+9xy-xy-3y^2)
Factor the 1st two and the last two terms separately to get:
=3(3x(x+3y)-y(x+3y))
Factor out the common factor to get:
=3(x+3y)(3x-y)
===========
Cheers,
Stan H.